OpenStudy (anonymous):

Factor the Polynomial 27x^3 – 27.

OpenStudy (solomonzelman):

\[27x^3=(3x)^3~~~~~~~~~~~27=3^3\]\[a^3-b^3=(a-b)(a^2+ab+b^2)\]